OpenStudy (anonymous):

please help A freight train completes its journey of 150 miles 1 hour earlier if its original speed is increased by 5 miles/hour. What is the train’s original speed? a. 20 miles/hour b. 25 miles/hour c. 30 miles/hour d. 35 miles/hour

OpenStudy (perl):

You can use the equation distance = rate * time

OpenStudy (perl):

A freight train completes its journey of 150 miles 1 hour earlier if its original speed is increased by 5 miles/hour. What is the train's original speed? We have two equations 150 = r*t 150 = ( r + 5) ( t - 1)

OpenStudy (anonymous):

how would I figure those out..? the r & t?

OpenStudy (perl):

you can solve this system by substitution. In the first equation solve for t in terms of r 150/r = t

OpenStudy (perl):

plug that into the second equation for t

OpenStudy (anonymous):

yeah... that confuses me...

OpenStudy (perl):

150 = r*t Solve for t t = 150/r substitute into second equation 150 = ( r + 5) ( 150/r - 1) distribute and simplify

OpenStudy (perl):

150 = r*150/r + 5*150/r - r - 5 150 = 150 + 5*150/r - r - 5 0 = 750 / r - r - 5

OpenStudy (perl):

multiply both sides by r 0 = 750 - r^2 - 5r

OpenStudy (perl):

then factor (or use quadratic formula)

OpenStudy (perl):

( r - 25 ) ( r + 30 ) = 0 r = 25 , - 30 we reject the negative solution